OpenStudy (anonymous):

for which equation is x=-1/3 NOT a solution? a. 7+8x=4 b. 6x-5=7 c. 18+9/x=-9 d. 12x+8=4

OpenStudy (anonymous):

7 + 8(-1/3) = 13/3 6(-1/3) - 5 = -7 18 + 9/(-1/3) = -9 12(-1/3) + 8 = 4 If answer b. actually says "6x - 5 = -7" then the answer is a.